kittygurl55:

5³ - 2 • 3 – (7² + 50)

Joe348:

First solve the problem in the parenthesis () \[7^2=49+50=99\] solve the exponents \[5^3=125\] \[125-2\times3-99\] \[20\]

Joe348:

Hope its understandable<3

kittygurl55:

thank u so much u made it really understandable

Joe348:

No problem glad to help<3